Ravi’s Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Ravi $5.55 per pound, and type B coffee costs $4.35 per pound. This month, Ravi made 125 pounds of the blend, for total cost of $626.55. How many pounds of type B coffee did he use?

Answer by ikleyn(52775)

.
Ravi’s Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Ravi $5.55 per pound,
and type B coffee costs $4.35 per pound. This month, Ravi made 125 pounds of the blend, for total cost of $626.55.
How many pounds of type B coffee did he use?
~~~~~~~~~~~~~~~~~


Let "b" be the pounds of coffee B (the unknown value under the problem's question).


Then the pounds of coffee A is  (125-b).


Write the total cost equation

    5.55(125-b) + 4.35b = 626.55.


Simplify and find "b"

    5.55*125 - 5.55b + 4.35b = 626.55

    -5.55b + 4.35b = 626.55 - 5.55*125

         -1.2b     =     -67.2

             b     =     -67.2/(-1.2) = 56.


ANSWER.  56 pounds of coffee B.

Solved.


---------------------


In this site, there is entire bunch of lessons covering various types of mixture problems.

Among these lessons,  the most relevant to your problem are the lessons
    - Word problems on mixtures for dry substances like coffee beans, nuts, cashew and peanuts
    - Word problems on mixtures for dry substances like candies, dried fruits
    - Word problems on mixtures for dry substances like soil and sand

Read them and become an expert in solution the mixture word problems.